OpenStudy (anonymous):

Find values for a and b in the equation ax^2-b=14 such that the solutions will be complex solutions.

OpenStudy (asnaseer):

ok, first write this in the standard form as:\[ax^2-bx-14=0\]I am assuming you missed the "x" from the b in your question above.

OpenStudy (anonymous):

Not necessarily lol. There is no x besides b.

OpenStudy (asnaseer):

ok, in that case we have:\[ax^2=14+b\]

OpenStudy (asnaseer):

which leads to:\[x^2=\frac{14+b}{a}\]for this to give complex solutions we need:\[\frac{14+b}{a}\lt0\]

OpenStudy (asnaseer):

because that would lead to generating a square root of a negative number which gives complex results

OpenStudy (anonymous):

I see now.. :) Thanks

OpenStudy (asnaseer):

you will need to find two inequalities here

OpenStudy (anonymous):

Indeeds.

OpenStudy (asnaseer):

ok - good - you DO understand :)

OpenStudy (anonymous):

Course! :)

OpenStudy (asnaseer):

great! it's just that it's difficult to judge how much each user already knows on a site like this.

OpenStudy (anonymous):

I understand :) Many levels of understanding in this site so it's a bit difficult.

OpenStudy (asnaseer):

but its great we have people like you who actually put in the effort of learning rather than just asking for answers.

OpenStudy (anonymous):

Oh I know, they just look for answers to input their answers and ask away next question.. I for one put in the effort to at least learn something, not just wasting peoples time :)

OpenStudy (asnaseer):

You will go far - wish you all the success in your life my friend.

OpenStudy (anonymous):

Much appreciated friend! I bid you the same.

OpenStudy (asnaseer):

thx - take care